# Impact of Weak Lensing Mass Mapping Algorithms on Cosmology Inference

### v1

@@ Abstract
Context. Weak gravitational lensing is a powerful tool for probing the distribution of dark matter in the Universe. Mass mapping algorithms, which reconstruct the convergence field from galaxy shear measurements, play a crucial role in extracting higher-order statistics from weak lensing data to constrain cosmological parameters. However, only limited research has been done on whether the choice of mass mapping algorithm affects the inference of cosmological parameters from weak lensing higher-order statistics.
Aims. This study aims to evaluate the impact of different mass mapping algorithms on the inference of cosmological parameters measured with weak lensing peak counts.
Methods. We employ Kaiser-Squires, inpainting Kaiser-Squires, and MCALens mass mapping algorithms to reconstruct the convergence field from simulated weak lensing data, generated from cosmo-SLICS simulations. Using these maps, we compute the peak counts and wavelet peak counts as our data vectors. We perform Bayesian analysis with MCMC sampling to estimate posterior distributions of cosmological parameters, including matter density, amplitude of matter fluctuations, and dark energy equation of state parameter.
Results. Our results indicate that the choice of mass mapping algorithm significantly affects the constraints on cosmological parameters, with the MCALens method improving constraints by up to 157% compared to the standard Kaiser-Squires method. This improvement arises from MCALens’s ability to better capture small-scale structures. In contrast, inpainting Kaiser-Squires yields constraints similar to Kaiser-Squires, indicating limited benefit from inpainting for cosmological parameter estimation with peaks.
Conclusions. The accuracy of mass mapping algorithms is critical for cosmological inference from weak lensing data. Advanced algorithms like MCALens, which offer superior reconstruction of the convergence field, can substantially enhance the precision of cosmological parameter estimates. These findings underscore the importance of selecting appropriate mass mapping techniques in weak lensing studies to fully exploit the potential of higher-order statistics for cosmological research.
@@ 1 Introduction
Weak gravitational lensing, or the distortion of distant galaxy images due to intervening mass along the line of sight, is a powerful cosmological probe (Kilbinger, 2015).
@@
In the past decade, weak lensing surveys, namely the Kilo Degree Survey (KiDS; Kuijken et al., 2019), the Dark Energy Survey (DES; Flaugher, 2005; DES Collaboration, 2022), and the Hyper Suprime Camera Survey (HSC; Aihara et al., 2017), and the Ultraviolet Near Infrared Optical Northern Survey (UNIONS, Guinot et al., 2022; Li et al., 2024), have imaged large areas of the extra-galactic sky to study the nature of dark matter, dark energy, and the sum of neutrino masses. The launch of the Euclid satellite (Euclid Collaboration et al., 2024) heralds a new generation of so-called Stage-IV weak lensing surveys, which will soon include the Vera C. Rubin Legacy Survey of Space and Time (LSST; LSST Dark Energy Science Collaboration, 2012), and the Nancy Grace Roman telescope (Spergel et al., 2015). These instruments will conduct weak lensing measurements over an unprecedented area and depth.
@@
Weak lensing surveys extract cosmological information primarily by measuring two-point summary statistics: the two-point galaxy shear correlation function and its Fourier equivalent, the power spectrum. However, current and future surveys are probing small angular scales with increasing precision. At these scales, weak lensing measurements contain non-Gaussian information due to non-linear structure formation that two-point statistics fail to capture (Weinberg et al., 2013). This has motivated the development of new summary statistics, known as higher-order statistics (HOS), designed to exploit the non-Gaussian information encoded in the matter density field. Some examples include peak counts (Marian et al., 2009; Dietrich & Hartlap, 2010; Lin & Kilbinger, 2015a, b; Harnois-Déraps et al., 2021; Ayçoberry et al., 2023; Harnois-Deraps et al., 2024), Minkowski functionals (Mecke et al., 1994; Kratochvil et al., 2012; Grewal et al., 2022), Betti numbers (Feldbrugge et al., 2019; Parroni et al., 2021), scattering transform coefficients (Cheng et al., 2020; Cheng & Ménard, 2021), three-point statistics and the bispectrum (Takada & Jain, 2003, 2004; Semboloni et al., 2011; Fu et al., 2014), map-level inference (Porqueres et al., 2022; Boruah & Rozo, 2024), and the starlet-$\ell_{1}$ norm (Ajani et al., 2021, 2023). Euclid Collaboration et al. (2023) have performed forecasts of ten different HOS with a unified framework and found that they all outperform two-point statistics in terms of constraining power on the cosmological matter density parameter, $\Omega_{\textrm{m}}$, the amplitude of the matter power spectrum, $\sigma_{8}$, and the dark energy equation of state parameter, $w_{0}$. While forecasts show the very promising potential constraining power of HOS, additional studies are required to demonstrate that systematic measurement errors can be controlled for these observables in practice.
@@
Several HOS are computed from the two-dimensional convergence, $\kappa$, which is not directly observable. Instead, $\kappa$ must be reconstructed from galaxy shears, in a process called ”mass-mapping”. This reconstruction is an ill-posed inverse problem, first because of galaxy shape noise and second because of missing data from observational artifacts such as bright stars, cosmic rays, and CCD defects. As a result, the fidelity of map reconstruction differs depending on the method used; see e.g. Jeffrey et al. (2021).
@@
In the literature, metrics to compare mass-mapping methods are typically based on the quality of the map reconstruction, with the implicit assumption that methods with a higher fidelity produce tighter cosmological contours. However, the effects of mass-mapping methods on cosmological parameters measured with HOS have not yet been fully quantified. Furthermore, Grewal et al. (2024) found that different choices in the inversion method alter cosmological contours in a Fisher-forecast analysis using Minkowski functionals.
In this work, we go a step further and measure the effect of different mass-mapping methods on cosmological constraints from peak counts in a full inference pipeline.
@@
Here, we focus on the following mass mapping methods: Kaiser-Squires (KS; Kaiser & Squires, 1993), inpainting Kaiser-Squires (iKS; Pires et al., 2009), and MCALens (Starck et al., 2021). KS is a linear method that is fast but implements no denoising, and does not account for missing data or boundary effects; iKS is an improvement upon KS that provides an iteratively applied inpainting scheme to reduce the effect of missing data. Finally, MCALens combines a sparsity-based method that captures non-Gaussian information on small scales with a Gaussian random field to replicate Gaussian behavior on large scales (Starck et al., 2021).
@@
We examine the effects of the preceding inversion methods on peak counts and wavelet peak counts. These summary statistics offer several unique advantages. They are fast to compute and have been shown to break the degeneracy between the standard model and fifth forces in the dark sector (Peel et al., 2018). Wavelet peak counts keep the advantages of peak counts but carry additional constraining power. In Stage-IV survey forecasts, Ajani et al. (2020) find that multi-scale wavelet peak counts yield tighter parameter constraints than a power spectrum analysis. This method is in fact so powerful that adding a power spectrum analysis does not further improve constraints compared to wavelet peak counts alone. Moreover, wavelet peak counts conveniently exhibit a covariance matrix that is nearly diagonal.
@@
The structure of this paper is as follows: in Sect. 2, we describe the SLICS simulations and their associated mock data, as developed in Harnois-Déraps et al. (2021), for use in our inference pipeline. We review the HOS that we measure (peaks, wavelet peaks) in Sect. 3, followed by a description of our cosmological inference pipeline in Sect. 4. We then compare the cosmological contours derived using each mass-mapping reconstruction in Sect. 5. Finally, we conclude the paper in Sect. 6.
@@ 2 Simulations
One method to compute a theoretical prediction of HOS as a function of cosmological parameters is to forward-model the HOS from cosmological simulations. In the following, we introduce the SLICS $N$-body simulations (Harnois-Déraps et al., 2019), and describe how they are used to construct mock galaxy catalogues of the DES Y1 footprint (Harnois-Déraps et al., 2021). We subsequently employ these mock galaxy catalogues to apply our mass-mapping algorithms and compute the HOS.
@@ 2.1 Cosmology Training Set
In this work, we use the version of cosmo-SLICS described in Harnois-Déraps et al. (2021). These simulations are designed for the HOS analysis of weak lensing data, covering points over a wide range of the $\left[\Omega_{\textrm{m}},\sigma_{8},h,w_{0}\right]$ parameter space, where $h$ is the Hubble parameter. They sample this space at 25 $w$CDM points as well as one $\Lambda$CDM point, arranged in a Latin Hypercube. At each of these points, a pair of $N$-body simulations with $1,536^{3}$ particles in a $(505\,h^{-1}$ Mpc $)^{3}$ co-moving volume is run, providing 2D projections of the matter density field. The precise mass of each particle is determined by the simulation volume and matter density and, therefore, varies with $h$ and $\Omega_{\textrm{m}}$, covering the range from $1.42$ to $7.63\times 10^{9}M_{\sun}$. The 2D projections are afterward organized into 10 (5 for each of the 2 random seeds) past light-cone density maps, $\delta_{\rm 2D}(\@vec{\theta},z)$, with a field of view of $100\,\mathrm{deg}^{2}$ at different redshifts $z$. Those maps are in turn used to obtain convergence and shear maps, $\kappa(\bm{\theta},z_{\textrm{s}})$ and $\gamma(\bm{\theta},z_{\textrm{s}})$, for a number of source redshifts $z_{\textrm{s}}$, as described in Section 2.3. The mean source redshift in each tomographic bin is $\langle z_{\textrm{s}}\rangle=0.403,0.560,0.773,0.984$.
@@ 2.2 Covariance Mocks
The covariance matrix, which expresses the sample variance of our HOS, is evaluated from the SLICS suite of simulations, described in Harnois-Déraps et al. (2018). This set is comprised of 124 fully independent $N$-body simulations, run at a fixed cosmology [^1] The cosmology of the Covariance Training Set adopts the parameters: $\Omega_{\textrm{m}}=0.2905$, $\Omega_{\Lambda}=0.7095$, $\sigma_{8}=0.826$, $h=0.6898$, and $n_{\textrm{s}}=0.969$. , but with different random phases in the initial conditions. The specifications of this set (volume, number of particles, resolution) are the same as those of the cosmology training set. We further increase the effective number of simulations by a factor of 10, by generating 10 different shape noise realizations for each of the 124 simulations. The covariance matrix is then computed from the $1,240$ realizations of this Covariance Training Set.
@@ 2.3 Matching Survey Properties
Harnois-Déraps et al. (2021) adapted the SLICS simulations to the DES Y1 survey by overlaying the simulated light cones with the data in a way that the resulting mock surveys have the $(\kappa,\gamma)$ values from the simulations at the positions and intrinsic ellipticities of the galaxies in the data. This was implemented by segmenting the DES Y1 catalogues into 19 tiles, each with a size of 100 $\rm deg^{2}$, and populating every simulated map with the galaxy samples from the corresponding tile. In particular, the positions of the individual observed galaxies were replicated in every simulated survey realization. Every simulated galaxy was assigned a redshift by sampling from the redshift distribution $n(z)$ corresponding to its tomographic bin.
In summary, each simulated light cone from the cosmology and covariance training sets is replicated 19 times, and linked to a full survey realization. As a final product, for the cosmology training set, we have 10 mock galaxy catalogue realizations for each of the 19 tiles that cover the DES Y1 footprint, at each of the 25 $w$CDM cosmologies and the single $\Lambda$CDM cosmology. For the covariance training set, we have 124 fully independent mock galaxy catalogue realizations at the fixed cosmology, for each of the 19 tiles.
3 Higher-Order Statistics on Weak Lensing Mass Maps
3.1 Mass mapping
The weak-lensing mass mapping problem is the task of reconstructing the convergence field $\kappa$ from the measured shear field $\gamma$. This is possible because $\kappa$ and $\gamma$ are not independent, and can both be expressed as second-order derivatives of the lensing potential $\psi$:
@@ 3.1 Mass mapping
where $\bm{k}$ is the wave vector in Fourier space. This estimator of the convergence field, however, is only optimal in the case of zero noise and no missing (or masked) data. Moreover, Eq. (2) is only defined for $\bm{k}\neq 0$, which means that the mean $\kappa$ cannot be recovered from $\gamma$ alone. This problem is known as mass-sheet degeneracy.
@@
To address these issues, several mass-mapping algorithms have been proposed in the literature. These techniques are based on different assumptions and methods and have been shown to produce different results in terms of the fidelity of the reconstructed maps. In this work, we compare the performance of the following mass-mapping techniques: the Kaiser-Squires (KS) method, the iterative Kaiser-Squires with inpainting (iKS), and the MCALens method. We provide a detailed description of these methods in App. A.
@@
The original KS method is a direct linear inversion of the shear field to the convergence field, using the relation Eq. (2). Since this method is very fast, it has been widely used in weak lensing studies (e.g., Ayçoberry et al., 2023; Euclid Collaboration et al., 2023; Dutta et al., 2024; Jeffrey et al., 2025). However, its simplicity leads to strong artifacts in the reconstructed convergence due to boundary effects, creates leakage between the E- and B-modes of the convergence field, and is not optimal in the presence of noise. The iKS algorithm is a variation of the KS method that iteratively applies the KS inversion along with an inpainting scheme that reduces the effects of missing data. The inpainting scheme uses the MCA iterative thresholding algorithm, introduced in Starck et al. (2005b, 2005), to fill in the missing convergence field data. This algorithm has been shown to improve the reconstruction of the convergence field in the presence of masks. Finally, the MCALens method is a non-linear mass-mapping algorithm that combines a Gaussian component to capture large-scale behavior and a non-Gaussian component, constructed with sparsity priors, to recover the strong peaks in the convergence field at small scales. As a non-linear method, MCALens is computationally more expensive, but it provides a more accurate reconstruction of the convergence field, especially at small scales, where the non-Gaussian features of the matter distribution are more pronounced (Starck et al., 2021).
@@
While there has been recent development of deep machine learning methods for mass mapping, e.g. Jeffrey et al. (2020); Remy et al. (2023), here we focus on unsupervised methods and leave comparison with deep learning methods for future work. Supervised methods, while promising, introduce issues such as training dataset representativity (i.e., the generalization problem). Moreover, in terms of mean square error, deep learning techniques have been shown to achieve results similar to those of MCALens (Remy et al., 2023). Additionally, although all of the above methods can be applied to data on the sphere, we instead take the approach of tiling the sky into small regions where the flat-sky approximation holds to ease comparison with the SLICS simulations during parameter inference, which are given in Cartesian coordinates.

### v2

# Impact of weak-lensing mass-mapping algorithms on cosmology inference

@@ Abstract
Context. Weak gravitational lensing is a powerful tool for probing the distribution of dark matter in the Universe. Mass-mapping algorithms, which reconstruct the convergence field from galaxy shear measurements, play a crucial role in extracting higher-order statistics from weak-lensing data to constrain cosmological parameters. However, only limited research has been done on whether the choice of mass-mapping algorithm affects the inference of cosmological parameters from weak-lensing higher-order statistics.
Aims. This study aims to evaluate the impact of different mass-mapping algorithms on the inference of cosmological parameters measured with weak-lensing peak counts.
Methods. We employed Kaiser-Squires, inpainting Kaiser-Squires, and MCALens mass-mapping algorithms to reconstruct the convergence field from simulated weak-lensing data, generated from cosmo-SLICS simulations. Using these maps, we computed the peak counts and multi-scale wavelet peak counts as our data vectors. We performed Bayesian analysis with Markov chain Monte Carlo sampling to estimate the posterior distributions of cosmological parameters, including the matter density, amplitude of matter fluctuations, and dark energy equation of state parameter.
Results. Our results indicate that the choice of mass-mapping algorithm significantly affects the constraints on cosmological parameters, with the MCALens method improving constraints by up to 157% compared to the standard Kaiser-Squires method. This improvement arises from MCALens’s ability to better capture small-scale structures. In contrast, inpainting Kaiser-Squires yields constraints similar to Kaiser-Squires, indicating a limited benefit from inpainting for cosmological parameter estimation with peaks.
Conclusions. The accuracy of mass-mapping algorithms is critical for cosmological inference from weak-lensing data. Advanced algorithms like MCALens, which offer superior reconstruction of the convergence field, can substantially enhance the precision of cosmological parameter estimates. These findings underscore the importance of selecting appropriate mass-mapping techniques in weak-lensing studies to fully exploit the potential of higher-order statistics for cosmological research.
@@ 1 Introduction
Weak gravitational lensing, or the distortion of distant galaxy images due to intervening mass along the line of sight, is a powerful cosmological probe (Kilbinger 2015). In the past decade, weak-lensing surveys, namely the Kilo Degree Survey (KiDS; Kuijken et al. 2019), the Dark Energy Survey (DES; Flaugher 2005; DES Collaboration 2022), and the Hyper Suprime Camera Survey (HSC; Aihara et al. 2017), and the Ultraviolet Near Infrared Optical Northern Survey (UNIONS, Guinot et al. 2022; Li et al. 2024), have imaged large areas of the extra-galactic sky to study the nature of dark matter, dark energy, and the sum of neutrino masses. The launch of the Euclid satellite (Euclid Collaboration et al. 2024) heralds a new generation of so-called Stage-IV weak-lensing surveys, which will soon include the Vera C. Rubin Legacy Survey of Space and Time (LSST; LSST Dark Energy Science Collaboration 2012), and the Nancy Grace Roman telescope (Spergel et al. 2015). These instruments will conduct weak-lensing measurements over an unprecedented area and depth.
@@
Weak-lensing surveys extract cosmological information primarily by measuring two-point summary statistics: the two-point galaxy shear correlation function and its Fourier equivalent, the power spectrum. However, current and future surveys are probing small angular scales with increasing precision. At these scales, weak-lensing measurements contain non-Gaussian information due to non-linear structure formation that two-point statistics fail to capture (Weinberg et al. 2013). This has motivated the development of new summary statistics, known as higher-order statistics (HOS), designed to exploit the non-Gaussian information encoded in the matter density field. Some examples include peak counts (Marian et al. 2009; Dietrich & Hartlap 2010; Lin & Kilbinger 2015a, b; Harnois-Déraps et al. 2021; Ayçoberry et al. 2023; Harnois-Déraps et al. 2024), Minkowski functionals (Mecke et al. 1994; Kratochvil et al. 2012; Grewal et al. 2022), Betti numbers (Feldbrugge et al. 2019; Parroni et al. 2021), scattering transform coefficients (Cheng et al. 2020; Cheng & Ménard 2021), three-point statistics and the bispectrum (Takada & Jain 2003, 2004; Semboloni et al. 2011; Fu et al. 2014), map-level inference (Porqueres et al. 2022; Boruah & Rozo 2024), and the starlet-$\ell_{1}$ norm (Ajani et al. 2021, 2023). Euclid Collaboration et al. (2023) performed forecasts of ten different HOS with a unified framework and found that they all outperform two-point statistics in terms of constraining power on the cosmological matter density parameter, $\Omega_{\textrm{m}}$, the amplitude of the matter power spectrum, $\sigma_{8}$, and the dark energy equation of state parameter, $w_{0}$. While forecasts show the very promising potential constraining power of HOS, additional studies are required to demonstrate that systematic measurement errors can be controlled for these observables in practice.
@@
Several HOS are computed from the two-dimensional convergence, $\kappa$, which is not directly observable. Instead, $\kappa$ must be reconstructed from galaxy shears, in a process called ”mass-mapping”. This reconstruction is an ill-posed inverse problem, first because of galaxy shape noise and second because of missing data from observational artefacts such as bright stars, cosmic rays, and CCD defects. As a result, the fidelity of map reconstruction differs depending on the method used (see e.g. Jeffrey et al. 2021).
@@
In the literature, metrics for comparing mass-mapping methods are typically based on the quality of the map reconstruction, with the implicit assumption that methods with a higher fidelity produce tighter cosmological contours. However, the effects of mass-mapping methods on cosmological parameters measured with HOS have not yet been fully quantified. Furthermore, Grewal et al. (2024), using Minkowski functionals, found that different choices in the inversion method alter cosmological contours in a Fisher-forecast analysis.
@@
In this work we go a step further and measure the effect of different mass-mapping methods on cosmological constraints from peak counts in a full inference pipeline. We focused on the following mass-mapping methods: Kaiser-Squires (KS; Kaiser & Squires 1993), inpainting Kaiser-Squires (iKS; Pires et al. 2009), and MCALens (Starck et al. 2021). KS is a linear method that is fast but implements no de-noising, and does not account for missing data or boundary effects; iKS is an improvement upon KS that provides an iteratively applied inpainting scheme to reduce the effect of missing data. Finally, MCALens combines a sparsity-based method that captures non-Gaussian information on small scales with a Gaussian random field to replicate Gaussian behaviour on large scales (Starck et al. 2021).
@@
We examined the effects of these three inversion methods on peak counts and wavelet peak counts. These summary statistics offer several unique advantages. They are fast to compute and have been shown to break the degeneracy between the standard model and fifth forces in the dark sector (Peel et al. 2018). Wavelet peak counts also have the advantages of peak counts but offer additional constraining power. In Stage-IV survey forecasts, Ajani et al. (2020) find that multi-scale wavelet peak counts yield tighter parameter constraints than a power spectrum analysis. This method is in fact so powerful that adding a power spectrum analysis does not further improve constraints compared to wavelet peak counts alone. Moreover, wavelet peak counts conveniently exhibit a covariance matrix that is nearly diagonal.
@@
The structure of this paper is as follows: In Sect. 2 we describe the cosmo-SLICS simulations and their associated mock data, as developed in Harnois-Déraps et al. (2021), that are used our inference pipeline. We review the HOS that we measure (peaks and wavelet peaks) in Sect. 3, which is followed by a description of our cosmological inference pipeline in Sect. 4. We then compare the cosmological contours derived using each mass-mapping reconstruction in Sect. 5. Finally, we conclude the paper in Sect. 6.
@@ 2 Simulations
One method for computing a theoretical prediction of HOS as a function of cosmological parameters is to forward-model the HOS from cosmological simulations. In the following, we introduce the SLICS $N$-body simulations (Harnois-Déraps et al. 2019), and describe how they are used to construct mock galaxy catalogues of the DES Y1 footprint (Harnois-Déraps et al. 2021). We employed these mock galaxy catalogues to apply our mass-mapping algorithms and compute the HOS.
@@ 2.1 Cosmology training set
In this work we used the version of cosmo-SLICS described in Harnois-Déraps et al. (2021). These simulations are designed for the HOS analysis of weak-lensing data, covering points over a wide range of the $\left[\Omega_{\textrm{m}},\sigma_{8},h,w_{0}\right]$ parameter space, where $h$ is the Hubble parameter. They sample this space at 25 $w$ cold dark matter (CDM) points as well as one $\Lambda$CDM point, arranged in a Latin Hypercube. At each of these points, a pair of $N$-body simulations with $1,536^{3}$ particles in a $(505\,h^{-1}$ Mpc $)^{3}$ co-moving volume is run, providing 2D projections of the matter density field. The precise mass of each particle is determined by the simulation volume and matter density and, therefore, varies with $h$ and $\Omega_{\textrm{m}}$, covering the range from $1.42$ to $7.63\times 10^{9}M_{\sun}$. The 2D projections are afterward organised into 10 (5 for each of the 2 random seeds) past light-cone density maps, $\delta_{\rm 2D}(\@vec{\theta},z)$, with a field of view of $100\,\mathrm{deg}^{2}$ at different redshifts $z$. Those maps are in turn used to obtain convergence and shear maps, $\kappa(\bm{\theta},z_{\textrm{s}})$ and $\gamma(\bm{\theta},z_{\textrm{s}})$, for a number of source redshifts $z_{\textrm{s}}$, as described in Section 2.3. The mean source redshift in each tomographic bin is $\langle z_{\textrm{s}}\rangle=0.403,0.560,0.773,0.984$.
@@ 2.2 Covariance mocks
The covariance matrix, which expresses the sample variance of our HOS, is evaluated from the SLICS suite of simulations, described in Harnois-Déraps et al. (2018). This set comprises 124 fully independent $N$-body simulations, run at a fixed $\Lambda$CDM cosmology [^1] The cosmology of the covariance training set adopts the parameters: $\Omega_{\textrm{m}}=0.2905$, $\Omega_{\Lambda}=0.7095$, $\sigma_{8}=0.826$, $h=0.6898$, and $n_{\textrm{s}}=0.969$. , but with different random phases in the initial conditions. The specifications of this set (volume, number of particles, resolution) are the same as those of the cosmology training set. We further increases the effective number of simulations by a factor of 10, by generating 10 different shape noise realisations for each of the 124 simulations. The covariance matrix is then computed from the $1,240$ realisations of this covariance training set.
@@ 2.3 Matching survey properties
Harnois-Déraps et al. (2021) adapted the SLICS simulations to the DES Y1 survey by overlaying the simulated light cones with the data in a way that the resulting mock surveys have the $(\kappa,\gamma)$ values from the simulations at the positions and intrinsic ellipticities of the galaxies in the data. This was implemented by segmenting the DES Y1 catalogues into 19 tiles, each with a size of 100 $\rm deg^{2}$, and populating every simulated map with the galaxy samples from the corresponding tile. In particular, the positions of the individual observed galaxies were replicated in every simulated survey realisation. Every simulated galaxy was assigned a redshift by sampling from the redshift distribution $n(z)$ corresponding to its tomographic bin.
In summary, each simulated light cone from the cosmology and covariance training sets is replicated 19 times and linked to a full survey realisation. As a final product, for the cosmology training set we have 10 mock galaxy catalogue realisations for each of the 19 tiles that cover the DES Y1 footprint, at each of the 25 $w$CDM cosmologies and the single $\Lambda$CDM cosmology. For the covariance training set, we have 124 fully independent mock galaxy catalogue realisations at the fixed cosmology, for each of the 19 tiles.
3 Higher-order statistics on weak-lensing mass maps
3.1 Mass-mapping
The weak-lensing mass-mapping problem is the task of reconstructing the convergence field $\kappa$ from the measured shear field $\gamma$. This is possible because $\kappa$ and $\gamma$ are not independent, and can both be expressed as second-order derivatives of the lensing potential $\psi$:
@@ 3.1 Mass-mapping
where $\bm{k}$ is the wave vector in Fourier space. This estimator of the convergence field, however, is only optimal in the case of zero noise, missing (or masked) data, and periodic boundary conditions. Moreover, Eq. (2) is only defined for $\bm{k}\neq 0$, which means that the mean $\kappa$ cannot be recovered from $\gamma$ alone. This problem is known as mass-sheet degeneracy.
@@
To address these issues, several mass-mapping algorithms have been proposed in the literature. These techniques are based on different assumptions and methods and have been shown to produce different results in terms of the fidelity of the reconstructed maps. In this work we compared the performance of the following mass-mapping techniques: the KS method, the iterative iKS, and the MCALens method. We provide a detailed description of these methods in Appendix A.
@@
The original KS method is a direct linear inversion of the shear field to the convergence field, using the relation Eq. (2). Since this method is very fast, it has been widely used in weak-lensing studies (e.g. Ayçoberry et al. 2023; Euclid Collaboration et al. 2023; Dutta et al. 2024; Jeffrey et al. 2025). However, its simplicity leads to strong artefacts in the reconstructed convergence due to boundary effects, creates leakage between the E and B modes of the convergence field, and is not optimal in the presence of noise. The iKS algorithm is a variation of the KS method that iteratively applies the KS inversion along with an inpainting scheme that reduces the effects of missing data. The inpainting scheme uses the morphological component analysis (MCA) iterative thresholding algorithm, introduced in Starck et al. (2005b, 2005), to fill in the missing convergence field data. This algorithm has been shown to improve the reconstruction of the convergence field in the presence of masks. Finally, the MCALens method is a non-linear mass-mapping algorithm that combines a Gaussian component to capture large-scale behaviour and a non-Gaussian component, constructed with sparsity priors, to recover the strong peaks in the convergence field at small scales. As a non-linear method, MCALens is computationally more expensive, but it provides a more accurate reconstruction of the convergence field, especially at small scales, where the non-Gaussian features of the matter distribution are more pronounced (Starck et al. 2021).
@@
While there has been recent development of deep machine learning methods for mass-mapping (e.g. Jeffrey et al. 2020; Remy et al. 2023), here we focus on unsupervised methods and leave comparison with deep learning methods for future work. Supervised methods, while promising, introduce issues such as training dataset representativity (i.e. the generalisation problem). Moreover, in terms of mean square error, deep learning techniques have been shown to achieve results similar to those of MCALens (Remy et al. 2023). Additionally, although all of the above methods can be applied to data on the sphere, we instead took the approach of tiling the sky into small regions where the flat-sky approximation holds to ease comparison with the SLICS simulations during parameter inference, which are given in Cartesian coordinates.